OpenStudy (anonymous):

what is the limit of sqrt(6-x) -2 / sqrt(3-x) -1 as x approaches 2?

OpenStudy (anonymous):

\[\lim \frac{ \sqrt{6-x} -2 }{ \sqrt{3-x}-1 }\]

OpenStudy (anonymous):

as x -> 2

OpenStudy (anonymous):

does it mean x=2

OpenStudy (anonymous):

well, it means as x approaches 2, so no then

OpenStudy (anonymous):

Multiply by the conjugates: \[\frac{ \sqrt{6-x} -2 }{ \sqrt{3-x}-1 }\cdot\frac{\sqrt{6-x} +2}{\sqrt{6-x} +2}\cdot\frac{\sqrt{3-x}+1}{\sqrt{3-x}+1}\]

OpenStudy (anonymous):

okay, I did that

OpenStudy (anonymous):

Right, so you have \[\frac{ 6-x -4 }{ 3-x-1 }\cdot\frac{{\sqrt{3-x}+1}}{\sqrt{6-x} +2}\] \[\frac{ 2-x }{ 2-x }\cdot\frac{{\sqrt{3-x}+1}}{\sqrt{6-x} +2}\] \[\frac{{\sqrt{3-x}+1}}{\sqrt{6-x} +2}\] Now directly substituting should work.

OpenStudy (anonymous):

ya, so i crossed out the first part of the denominator with the first part of the numerator

OpenStudy (anonymous):

and then I became stuck

OpenStudy (anonymous):

do I just plug 2 into x now is my question

OpenStudy (anonymous):

Yeah, that's what I meant by directly substituting.

OpenStudy (anonymous):

okay, ya. So i was on the right track, but I doubted myself. Thank you so much :)

OpenStudy (anonymous):

You're welcome